### v3

# Sharp Bounds for Multiple Models in Matrix Completion

Abstract
In this paper, we demonstrate how a class of advanced matrix concentration inequalities, introduced in Brailovskaya and van Handel [2024], can be used to eliminate the dimensional factor in the convergence rate of matrix completion. This dimensional factor represents a significant gap between the upper bound and the minimax lower bound, especially in high dimension. Through a more precise spectral norm analysis, we remove the dimensional factors for three popular matrix completion estimators, thereby establishing their minimax rate optimality.
1 Introduction
Matrix recovery from a small subset of entries, often referred to as matrix completion, is a classical problem in high-dimensional statistics with wide-ranging applications across various fields. Numerous frameworks and methods have been developed for matrix completion, with a fundamental understanding that a large matrix cannot be recovered unless it exhibits certain low-rank properties. As a convex relaxation of rank penalization, nuclear norm penalized trace regression has become a popular approach and corresponding theoretical research has been fruitful. See, for example, Candès and Tao [2010], Recht [2011], Koltchinskii et al. [2011], Negahban and Wainwright [2012], Klopp [2014], Klopp et al. [2017], Minsker [2018], Yu et al. [2024].
In particular, many studies focus on the convergence rate of the nuclear norm penalization methods Koltchinskii et al. [2011], Negahban and Wainwright [2012], Klopp [2014], Yu et al. [2024]. However, the minimax lower bound, obtained in Koltchinskii et al. [2011] and Negahban and Wainwright [2012], does not include a logarithmic dimension factor that always appears in the upper bound of the convergence rate. This discrepancy between the two bounds becomes more pronounced in high-dimensional settings, highlighting a potential gap in the existing theoretical understanding. Due to the ubiquitousness of this issue, all the previous works have had to qualify their results by stating “our estimator is minimax optimal, up to a logarithmic factor”. In this paper, we leverage a class of sharp matrix concentration inequalities to remove the logarithmic dimension factor. After our work, the qualification will no longer be necessary.
Problem setup
Let $A_{0}\in\mathbb{R}^{m_{1}\times m_{2}}$ be the unknown matrix we aim to recover. We observe independent samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$ from the model
$Y_{i}=\langle X_{i},A_{0}\rangle+\xi_{i},\quad i=1,\ldots,n,$ (1)
where $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$, $\langle X_{i},A_{0}\rangle=\text{tr}(X_{i}^{T}A_{0})$, and $\text{tr}(B)$ denotes the trace of matrix $B$. The sampling matrices $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ take values from $\{e_{j}(m_{1})e_{k}^{T}(m_{2}),1\leq j\leq m_{1},1\leq k\leq m_{2}\}$, where $\{e_{j}(m_{1})\}_{j=1}^{m_{1}},\{e_{k}(m_{2})\}_{k=1}^{m_{2}}$ denote standard basis vectors in $\mathbb{R}^{m_{1}},\mathbb{R}^{m_{2}}$, respectively. This implies that each $X_{i}$ only has one non-zero entry, indicating the location of the sampled entry.
The distribution of the noise variables $\xi_{i}$ will be specified in different settings. The sampling distribution of $X_{i}$ is denoted by
$P_{jk}=\mathbb{P}(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2})),\quad 1\leq j\leq m_{1},1\leq k\leq m_{2}.$
$X_{i}$ and $\xi_{i}$ can be dependent. Throughout the paper, we consider that $\text{rank}(A_{0})\leq r$ and $\|A_{0}\|_{\infty}\leq a$, where $\|A\|_{\infty}=\max_{1\leq j\leq m_{1},1\leq k\leq m_{2}}|a_{jk}|$ is the entrywise maximum norm of $A=(a_{jk})$.
The logarithmic dimension factor issue
The matrix completion problem is often addressed by a penalized regression. Assuming $A_{0}$ is of low rank, a natural choice for the penalty is the rank of a matrix. However, the rank penalty is non-convex and makes the corresponding program difficult to solve. The nuclear norm $\|\cdot\|_{*}$, namely the sum of all the singular values of a matrix, is usually used as a convex relaxation for the rank. There are many classical works on nuclear norm penalization methods under different assumptions. Assuming the noise has exponential decay, Koltchinskii et al. [2011] consider the case that the sampling distribution is known; Negahban and Wainwright [2012] study the situation where the row indices and the column indices of the observations are sampled independently and the penalty is a weighte nuclear norm; Klopp [2014] explores the problem for general sampling distributions. Consider the classical nuclear norm penalized least squares estimator
$\hat{A}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{*}\right\}.$
Assuming $\text{rank}(A_{0})\leq r$, $X_{i}$ has a general non-uniform distribution and the noise has exponential decay, Klopp [2014] obtains the following bound that holds with high probability:
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprle\log(m_{1}+m_{2})\dfrac{r\max(m_{1},m_{2})}{n},$ (2)
[…]
$\inf_{\hat{A}}\sup_{\mathcal{A}}\mathbb{P}\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\dfrac{r\max(m_{1},m_{2})}{n}\right)\geq C,$
where $C\in(0,1)$ is a constant, and the infimum and supremum are taken over all the possible estimators and the parameter space $\{A_{0}\in\mathbb{R}^{m_{1}\times m_{2}}:\text{rank}(A_{0})\leq r,\|A_{0}\|_{\infty}\leq a\}$ respectively. The logarithmic dimension factor $\log(m_{1}+m_{2})$, as the gap between the upper and lower bounds, exists in other matrix completion settings as well. See, for instance, Klopp et al. [2017] for corrupted matrix completion, Yu et al. [2024] for matrix completion with heavy tailed noise.
By our analysis, this gap arises essentially from the spectral norm analysis of some random matrices. In matrix completion, we frequently need to control the spectral norms of random matrices of the form $\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}X_{i}$, where $\zeta_{i}$ are bounded or light-tailed random variables. The use of traditional concentration inequalities in this context typically introduces a logarithmic dependence on the dimension. This limitation is not unique to matrix completion—it arises in a variety of statistical and mathematical problems and has garnered significant attention in recent years. Notably, Bandeira et al. [2023] and Brailovskaya and van Handel [2024] proposed sharp new concentration inequalities that can successfully remove the logarithmic dependence in various settings. These results will serve as key tools in our paper.
Another point worth mentioning is that there is an alternative sampling model in matrix completion. In this model, there are Bernoulli random variables $\eta_{ij}$ indicating whether the entry of $A_{0}$ at location $(i,j)$ is observed. Under this assumption, the observations can be written as:
[…]
where $\xi_{ij}$ represents the noise at location $(i,j)$ and $A_{0}=(a_{0,ij})$. This sampling model is often referred to as sampling without replacement because each entry can be selected at most once. A rich body of literature has investigated this model [Candes and Plan, 2010, Chen and Wainwright, 2015, Candes and Recht, 2012, Keshavan et al., 2009, Chen and Wainwright, 2015, Chen et al., 2020]. As a contrast, the sampling strategy considered in our paper is known as sampling with replacement, as it allows repeated observations of the same entry. Since $n$ is often assumed to be much smaller than the matrix dimension $m_{1}\times m_{2}$, the number of repeated observations could be very small. Despite different sampling strategies, the theoretical results derived in both models are often compared side by side [Chen and Wainwright, 2015, Chen et al., 2020]. For instance, some works derive bounds of the following form [Chen and Wainwright, 2015]:
$\quad\frac{\|\widehat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\frac{r\max(m_{1},m_{2})}{m_{1}m_{2}p},$ (3)
under the non-replacement model with $\eta_{ij}\sim\text{Bernoulli}(p)$. Here, as the expected sample size, $m_{1}m_{2}p$ plays a role as $n$ in the with-replacement model. Compared to (3), the extra $\log(m_{1}+m_{2})$ factor in (2) is a common critique of algorithms developed under the with-replacement model[Chen and Wainwright, 2015, Chen et al., 2020]. Therefore, removing this logarithmic factor can help significantly defend the theoretical validity of the algorithms [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014].
Our contribution
In this paper, we revisit three estimators that have been studied in other research works:
  • The unknown matrix is of low rank and the noise is heavy tailed. The estimator in Yu et al. [2024] is designed for the case that the noise is only assumed to have finite second moment.
  • The unknown matrix is of low rank and the noise is sub-Gaussian. Nuclear norm penalized least squares is a common approach under this setting [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014]. We consider the estimator proposed in Klopp [2014] to illustrate our new results.
  • The unknown matrix is of low rank and the noise is sub-Gaussian with unknown variance. The parameter tuning in matrix completion often requires prior knowledge of the noise variance. Klopp [2014] proposes an estimator to solve such a problem.
All the previous results on the convergence rates of the above three estimators involve the logarithmic dimension factor. We remove the dimensional factor for all the three estimators, thereby proving their minimax rate optimality.
The remainder of the paper is organized as follows. Continuing in this section, we introduce the notation. In Section 2, we present our main results. We give some concluding remarks in Section 3. Section 4 is dedicated to the proofs.
Notation
To present our results, we collect the notations used throughout this paper for reader’s convenience.
  • $M=m_{1}\vee m_{2}$, $m=m_{1}\wedge m_{2}$, $d=m_{1}+m_{2}$.
  • Let $A=(a_{ij})$, for $1\leq i\leq m_{1},1\leq j\leq m_{2}$. Let $\{\sigma_{l}\}_{l=1}^{m}$ denote ordered singular values of $A$: $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}$. $\|A\|_{F}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{2}}a_{ij}^{2}}$, $\|A\|_{*}=\sum_{l=1}^{m}\sigma_{l}$, $\|A\|_{\infty}=\max_{i\leq m_{1},j\leq m_{2}}|a_{ij}|$, $\|A\|=\sigma_{1}$. $\|A\|_{w(F)}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{i=1}^{m_{2}}a^{2}_{ij}P_{ij}}$ is a weighted Frobenius norm associated with the distribution of the sampling matrices $X_{i}$.
  • $C,C_{1},C_{2},\ldots$ are positive constants whose value may vary at each occurrence.
  • For a matrix $A$, denote its singular value decomposition as $A=U\Sigma V^{T}$. We define the projection operators
    $P_{A}(B)=UU^{T}BVV^{T}$
    and
    $P_{A}^{\perp}(B)=B-P_{A}(B).$
  • $\{\epsilon_{i}\}_{i=1}^{n}$ are i.i.d. symmetric Bernoulli random variables, namely
    $\mathbb{P}(\epsilon_{i}=1)=\mathbb{P}(\epsilon_{i}=-1)=\frac{1}{2}.$
    They are independent of the observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$.
  • We write $(\log d)^{\alpha}$ as $\log^{\alpha}d$ and $(\mathbb{E}[X])^{\alpha}$ as $\mathbb{E}^{\alpha}[X]$ for aesthetic reasons.
  • A random variable $X$ is called $\psi_{\alpha}$ random variable if it has finite $\psi_{\alpha}$ norm,
    $\|X\|_{\psi_{\alpha}}=\inf\left\{C>0;\mathbb{E}\exp\left(\left(\frac{|X|}{C}\right)^{\alpha}\right)\leq 2\right\}.$
    In particular, when $\alpha=2$, $\|\cdot\|_{\psi_{2}}$ is called sub-gaussian norm and X is called sub-gaussian random variable.
2 Main Results
In this section, we present new upper bounds for the three estimators mentioned in Section 1. In particular, compared to the existing bounds, the new bounds do not involve the dimension factor $\log d$ and hence match with the minimax lower bounds.
At first, we introduce assumptions used in our paper. Regarding the noise $\xi_{i}$, they can be heavy tailed or sub-Gaussian in different contexts and their distributions can depend on the sampled entry. We introduce the following assumptions.
Assumption 1.
Given $X_{i}$, the noise variables satisfy
$E[\xi_{i}|X_{i}]=0$
and their second moments are uniformly bounded
$E[\xi_{i}^{2}|X_{i}]\leq\sigma^{2}.$
Assumption 2.
The noise variables $\xi_{i}$ satisfy $E[\xi_{i}|X_{i}]=0$, $E[\xi^{2}_{i}]=\sigma^{2}$ and given $X_{i}$, it holds uniformly
$E\exp(t\xi_{i}/\sigma)\leq\exp(L_{1}^{2}t^{2}),\forall t\in\mathbb{R},$
where $L_{1}>0$ is a constant.
Regarding the distribution of $X_{i}$, we consider a non-uniform sampling distribution.
Assumption 3.
$X_{i}$’s distribution satisfies
$(1)\leavevmode\nobreak$ $\max_{j}\sum_{k=1}^{m_{2}}P_{jk}\leq\frac{L_{2}}{m},$
$(2)\leavevmode\nobreak$ $\max_{k}\sum_{j=1}^{m_{1}}P_{jk}\leq\frac{L_{2}}{m},$
$(3)\leavevmode\nobreak$ $\min_{j,k}P_{jk}\geq\frac{1}{\mu m_{1}m_{2}},$
$(4)\leavevmode\nobreak$ $\max_{j,k}P_{jk}\leq\frac{L_{3}}{m\log^{3}d},$
where $L_{2},L_{3}\geq 0$ are constants, and $\mu\geq 1$ is allowed to change with the sample size or dimension.
Conditions $(1)$-$(3)$ of Assumption 3 are the same as in Klopp [2014]. $L_{2}$ is introduced to control the row and column sampling probabilities, preventing them from being too large. Condition (4) requires that the sampling doe not extremely concentrate on a few entries. Note that Condition (1) or (2) already implies $\max_{j,k}P_{jk}\apprle\frac{1}{m}$ which only differs from Condition (4) by a logarithmic factor. Hence, (4) is a very mild condition.
In following subsections, we introduce our $\log d$-free error bounds.
2.1 Matrix completion with heavy tailed noise
We start with our improvement on a recent advance in matrix completion [Yu et al., 2024]. Heavy-tailed data is ubiquitous in many fields such as finance, microeconomics and biology. To address heavy-tailed noise in the context of matrix completion, several estimators have been proposed. Assuming that the second moment of the noise is finite, Minsker [2018] proposes a two-step procedure, where a truncated-type estimator $\tilde{A}$ is obtained first and then a nuclear norm penalized regression is performed on $\tilde{A}$. Fan et al. [2021] also proposes a two-step procedure while a stronger moment condition $E\xi_{i}^{2k}<\infty$ for $k>1$ is required. Elsener and van de Geer [2018] considers a class of loss functions including Huber loss function. They assume the distribution of the noise is symmetric and some additional conditions hold. In particular, Yu et al. [2024] only assumes that the noise variables have finite second moments and still derives strong properties for the estimator under study. The estimator proposed in their paper is based on the Huber loss function, which is one of the fundamental tools for addressing heavy tailed noise.
We first define
$l_{\tau}(x)=\begin{cases}\frac{x^{2}}{2},\leavevmode\nobreak\ \leavevmode\nobreak\ |x|\leq\tau,\\ \tau|x|-\frac{\tau^{2}}{2},\leavevmode\nobreak\ \leavevmode\nobreak\ |x|>\tau,\\ \end{cases}$
and Huber loss is defined as $\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$. The estimator to be studied is
$\hat{A}_{H}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)+\lambda\|A\|_{*}\right\}.$ (4)
We obtain the following new convergence rate for $\hat{A}_{H}$.
Theorem 1.
Let Assumptions 1 and 3 hold. We additionally assume $n\geq 160m\log^{4}d$ and $\xi_{i}|X_{i}$ are symmetric. If we take
$\lambda=C_{1}\max(\sigma,a)\sqrt{\frac{1}{nm}},$
$\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}},$
with probability at least $1-\frac{2}{d}$, the following bound holds
$\dfrac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\leq C_{2}\frac{\max(a^{2},\sigma^{2})rM}{n}.$ (5)
Here, $C_{1}$ and $C_{2}$ depend on the constants $L_{2},L_{3}$ from Assumption 3.
Yu et al. [2024] provides the following theorem.
Theorem 2.
Let Assumption 1 hold and set $P_{jk}=\frac{1}{m_{1}m_{2}}$. If we take
$\lambda=C_{1}\max(a,\sigma)\sqrt{\frac{\log d}{nm}},$
$\tau=C_{2}\max(a,\sigma)\sqrt{\frac{n}{m\log d}},$
then with probability at least $1-\frac{3}{d}$, we have
$\dfrac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\max(a^{2},\sigma^{2})\dfrac{rM\log d}{n},a^{2}\sqrt{\frac{\log d}{n}}\right\}$
provided that $n\geq C_{4}m\log d$. Here, $C_{1}\sim C_{3}$ are universal constants.
Theorem 1 improves over Theorem 2 in several aspects.
  1. 1.
    Our new result eliminates the $\log d$ factor in the convergence rate. It is achieved by utilizing new advanced concentration inequalities in Brailovskaya and van Handel [2024]. See our Lemma 14 for a detailed analysis. Koltchinskii et al. [2011] provides the following minimax lower bound in the case of gaussian noise and uniform sampling distribution:
    $\inf_{\hat{A}}\sup_{A_{0}}\mathbb{P}\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\geq C_{1}\dfrac{\min(\sigma^{2},a^{2})rM}{n}\right)\geq C_{2},$ (6)
    where $C_{1},C_{2}$ are universal constants, and the infimum and supremum are taken over all possible estimators and rank-$r$ matrices respectively. Combining the new bound (5) with the minimax lower bound (6) shows that the removal of the dimension factor $\log d$ establishes the minimax rate optimality of the estimator (4).
  2. 2.
    Theorem 1 allows for a more general sampling distribution while Theorem 2 requires uniform distribution. We also eliminate the nuisance term $O(\sqrt{(\log d)/n})$ from the convergence rate, via a different peeling argument [Wainwright, 2019] in the analysis of certain empirical processes.
  3. 3.
    Our analysis reveals a more precise choice for the tuning parameter $\lambda$: the optimal choice for $\lambda$ should be of the order $O(\sqrt{1/(nm)})$ instead of $O(\sqrt{(\log d)/(nm)})$, which sheds light on the practice of parameter tuning.
  4. 4.
    Theorem 1 requires symmetry of the noise, while Theorem 2 does not. If the noise is not symmetric and the sampling distribution is uniform, the bound (5) continues to hold under the additional condition $M\geq m\log^{5}d$. While this condition is very mild, this seems to be a pure technical condition and is expected to be unnecessary under future investigation. Moreover, assuming that the $\xi_{i}$ have a finite moment of order $2+\kappa$ for some $\kappa>0$, a $\log d$-free bound can be established under a slightly weaker condition on sample size:
    $n\geq\max\left(m\log^{4+\frac{1}{\kappa+1}}d,\log^{\frac{4}{\kappa}}d\right).$
    See our analysis in the proof of inequality (5).
2.2 Matrix completion with known variance
The setting that the noise have distributions with exponential decay is most studied, see, e.g., Koltchinskii et al. [2011], Negahban and Wainwright [2012], Klopp [2014]. Under the assumption of light tailed noise, an estimator proposed in Klopp [2014] is:
$\hat{A}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{*}\right\}.$ (7)
We now present our new result:
Theorem 3.
Let Assumptions 2 and 3 hold. Assume that $n\geq C_{1}m\log^{4}d\leavevmode\nobreak\ (\log d+\log n)$. If we take
$\lambda=C_{2}\sigma\sqrt{\frac{1}{nm}},$
the following bound holds with probability at least $1-\frac{2}{d}$
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{3}\mu^{2}\max(a^{2},\sigma^{2})rM}{n}.$ (8)
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2},L_{3}$ from Assumptions 2-3.
To compare with existing bounds, we mention the following theorem:
Theorem 4 (Theorem 7 in Klopp [2014]).
Let Assumption 2, Conditions (1)-(3) in Assumption 3 hold. Assume that $n\geq C_{1}m\log^{2}d$. If we take
$\lambda=C_{2}\sigma\sqrt{\dfrac{\log d}{nm}},$
the following bound holds with probability $1-\frac{3}{d}$:
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},a^{2}\mu\sqrt{\frac{\log d}{n}}\right\}.$
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2}$ from Assumptions 2-3.
Comparing Theorems 3 and 4, we make several remarks:
  1. 1.
    Theorem 3 eliminates a $\log d$ factor in the convergence rate. Comparing (8) and (6), the removal of the $\log d$ factor establishes the minimax optimality of estimator (7).
  2. 2.
    Theorem 3 also provides the correct order of tuning parameter $\lambda$ for (7). The old one is $O(\sqrt{\log d/(nm)})$ while the correct one in Theorem 3 is $O(\sqrt{1/(nm)})$. The $\log d$ difference can be significant in high dimensional settings.
  3. 3.
    Theorem 4 involves an additional nuisance term $O(\sqrt{\log d/n})$, similarly to Theorem 2. In the high- dimensional regime where $n\leq m_{1}m_{2}$, it is straightforward to verify that the first term–the desirable term–is the larger one. However, in less challenging scenarios in which $n$ is sufficiently large, the nuisance term dominates. This limitation arises from a Frobenius norm based peeling argument that Klopp [2014] used to prove restricted strong convexity. In contrast, we have adopted a different peeling scheme introduced in Wainwright [2019], which peels certain matrix space by the infinity norm and the nuclear norm. As a result, our new bound does not include the nuisance term $O(\sqrt{\log d/n})$.
  4. 4.
    Our new bound requires a slightly stronger condition on $n$. The $\log n$ term arises from the truncation of the sub-Gaussian noise, in order to apply the matrix concentration inequalities from Brailovskaya and van Handel [2024]. In the most interesting regime where $n\leq m_{1}m_{2}$, the sample size requirement in Theorem 3 becomes $n\geq Cm\log^{5}d$ which is in the typical form of $n\geq m\text{Poly}(\log d)$ in the matrix completion literature. Under both with-replacement and without-replacement settings, the order of $\text{Poly}(\log d)$ often vary with problem assumptions and methods used; see, for example, Koltchinskii et al. [2011], Klopp [2014], Chen et al. [2020]. In our results, the leverage of the concentration inequalities [Brailovskaya and van Handel, 2024] causes a small change of the order of $\text{Poly}(\log d)$. This fluctuation is acceptable in light of the substantial gain in sharpness of the resulting bounds. We emphasize that previous concentration inequalities cannot yield bounds as sharp as ours, even when the sample size is fixed at the same level as in our results.
2.3 Matrix completion with unknown variance of the noise
The choice of $\lambda$ in Theorem 4 is dependent on the variance of the noise. In practice, $\sigma$ is often unknown. To deal with this issue, Klopp [2014] proposes the following square-root lasso type estimator and studies its convergence rate.
$\hat{A}_{S}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\sqrt{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}+\lambda\|A\|_{*}\right\}.$ (9)
We obtain a new result for (9).
Theorem 5.
Let Assumptions 2, 3 hold. Let $n\geq C_{1}m\log^{4}d\leavevmode\nobreak\ (\log d+\log n)$. We can take
$\lambda=C_{2}\sqrt{\frac{1}{nm}},$
and with probability at least $1-\frac{2}{d}-2\exp(-C_{3}n)$, the following bound holds
$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{4}\mu^{2}\max(a^{2},\sigma^{2})rM}{n},$ (10)
provided that $n\geq 16\mu Mr$. Here, $C_{1}$ is a universal constant, $C_{3}$ depends on $L_{1}$, and $C_{2},C_{4}$ depend on $L_{1},L_{2},L_{3}$, where $L_{i}$’s are constants from Assumptions 2-3.
We can compare the above bound with the one obtained in Klopp [2014].
Theorem 6 (Theorem 10 in Klopp [2014]).
Let Assumption 2 and and Conditions (1)-(3) in Assumption 3 hold. We can take
$\lambda=C_{1}\sqrt{\frac{\log d}{nm}},$
and the following upper bound holds with probability $1-\frac{3}{d}-2\exp(-C_{2}n)$,
$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},a^{2}\mu\sqrt{\frac{\log d}{n}}\right\},$
provided that $n\geq 8C_{1}\mu rM\log d$. Here $C_{1},C_{3}$ depend on $L_{1},L_{2}$ and $C_{2}$ depends on $L_{1}$, where $L_{1},L_{2}$ are constants from Assumptions 2-3.
Comparing the convergence rate in Theorem 6 and the lower bound in (6), they are not matched due to the existence of the $\log d$ factor. Our new bound in (10) resolves the issue and shows that the estimator (9) is indeed minimax rate optimal. Regarding more comparisons between Theorems 5 and 6, we refer to remarks after Theorem 4 for similar comments.
3 Conclusion
In this paper, we considered three common matrix completion settings: (1) the unknown matrix is of low rank and the noise is heavy tailed; (2) the unknown matrix is of low rank and the noise is sub-Gaussian with known variance; (3) the unknown matrix is of low rank and the noise is sub-Gaussian with unknown variance. We revisited three popular estimators in these settings and developed sharper upper bounds to establish their minimax rate optimality.
A key technical contribution lies in employing powerful matrix concentration inequalities introduced in Brailovskaya and van Handel [2024] to successfully remove the dimension factor $\log d$ that has remained in the upper bounds prior to our work. As previously noted, the $\log d$ factor is ubiquitous in matrix completion problems under the sampling with replacement setting. Our sharp analyses (e.g., Lemma 14 establishes dimension-free results for a variety of random matrices) can be applied or adapted to improve other existing results on matrix completion [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp et al., 2017, Chen et al., 2024].
4 Proof
Throughout this section, we denote
$L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle).$
The derivative of $l_{\tau}(\cdot)$ is denoted by $\phi_{\tau}(\cdot)$, where
$\phi_{\tau}(x)=\begin{cases}x,|x|\leq\tau,\\ \operatorname{sign}(x)\tau,|x|>\tau.\end{cases}$
In this section, $c_{0}\sim c_{4}$ are specific constants whose values do not change at each occurrence.
4.1 Proof of Theorem 1
We begin with the following general result.
Lemma 7.
Let Assumptions 1 and 3 hold. Suppose $n\geq 160m\log^{4}d$ and $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$. If we take
$\lambda\geq 3\left\|\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\right\|,$
then the following bound holds with probability at least $1-\frac{1}{d}$,
$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}rm_{1}m_{2}(\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|)+\frac{\mu^{2}\log d}{n}.$
Here $\mathcal{R}$ is defined as $\mathcal{R}=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}},$ where $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
Proof of Theorem 1.
Recall $\tau=\frac{\max(a,\sigma)}{\log^{2}d}\sqrt{\frac{n}{m}}$. By (30), we obtain with probability at least $1-\frac{1}{d}$,
$\left\|\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\right\|\leq C_{1}\max(\sigma,a)\sqrt{\dfrac{1}{nm}},$
where $C_{1}$ depends on $L_{2}\sim L_{3}$. With this specific $C_{1}$, one can take $\lambda=3C_{1}\max(\sigma,a)\sqrt{\frac{1}{nm}}$ to apply Lemma 7. By (32), we obtain
$\mathbb{E}\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|\leq C_{2}\sqrt{\frac{1}{nm}},$ (11)
where $C_{2}$ depends on $L_{2}\sim L_{3}$. Substituting this bound and the above choice of $\lambda$ into the inequality from Lemma 7, we complete the proof of Theorem 2. ∎
Remark.
Without the condition that $\xi_{i}$ are symmetric, $\nabla L(A_{0})$ may not have mean 0. In this case, it is necessary to control the spectral norm of its expectation. Specifically, we write:
$\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}=\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(\xi_{i})-\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}])X_{i}+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}]X_{i}.$
Denote $E[\phi_{\tau}(X_{i})|X_{i}]$ by $u_{i}$ (we omit the dependence on $\tau$ to ease the notation). While (30) still applies to $\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(\xi_{i})-u_{i})X_{i}$ given $\tau=\max(\sigma,a)\sqrt{\frac{n}{m\log^{4}d}}$, we need to estimate
$\left\|\frac{1}{n}\sum_{i=1}^{n}u_{i}X_{i}\right\|$ $=\left\|\dfrac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[u_{i}X_{i}]\right\|$
$\leq\left\|\dfrac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\right\|+\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[u_{i}X_{i}]\right\|.$
Assume a random variable $Y$ has 0 mean and finite moment of order $2+\kappa$ with $\mathbb{E}^{\frac{1}{2+\kappa}}[|Y|^{2+\kappa}]\leq\sigma$ where $\kappa\geq 0$. By Cauchy-Schwarz inequality, for a positive number $\tau$,
$|\mathbb{E}\phi_{\tau}(Y)|$ $=|\mathbb{E}[Y-\phi_{\tau}(Y)]|$
$=|\mathbb{E}[(Y-\phi_{\tau}(Y)\textbf{1}(|Y|>\tau)]$
$\leq\mathbb{E}^{1/p}[|Y|^{p}]\mathbb{P}^{1/q}(Y>\tau)$
where $11/p+1/q=$ 1. Note $\mathbb{P}(Y>\tau)\leq\mathbb{E}[|Y|^{2+\kappa}]/\tau^{2+\kappa}$. Take $1/q=1-1/(2+\kappa)$. Then, we obtain
$|\mathbb{E}[(Y-\phi_{\tau}(Y))\textbf{1}(|Y|>\tau)]|\leq\mathbb{E}^{\frac{1}{2+\kappa}}[|Y|^{2+\kappa}](\mathbb{E}^{\frac{1+\kappa}{2+\kappa}}[|Y|^{2+\kappa}]\tau^{-(1+\kappa)}).$
Let $\tau=\max(\sigma,a)\sqrt{\frac{n}{m\log^{4}d}}$. By the above inequality, we obtain $|u_{i}|\leq\max(\sigma,a)(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$. Assuming $P_{jk}=\frac{1}{m_{1}m_{2}}$, we apply Theorem 7.3.1 in Tropp et al. [2015] to $\frac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])$. Since $\frac{1}{n}\|u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}]\|$ is bounded by $\max(\sigma,a)\frac{2}{n}(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$, the following holds with probability at least $1-\frac{1}{d}$,
$\left\|\frac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\right\|\lesssim\max(\sigma,a)\left(\frac{m\log^{4}d}{n}\right)^{(1+\kappa)/2}\sqrt{\frac{\log d}{nm}}+\max(\sigma,a)\frac{\log d}{n}\left(\frac{m\log^{4}d}{n}\right)^{(1+\kappa)/2},$
where we use easy estimates $\|\mathbb{E}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])^{T}\|\lesssim\max(\sigma,a)\frac{L_{2}}{m}(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$ and $\|\mathbb{E}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])^{T}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\|\lesssim\max(\sigma,a)\frac{L_{2}}{m}(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$. Given $n\geq m\log^{4+1/(1+\kappa)}d$, we have
$\left\|\dfrac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\right\|\lesssim\sigma\sqrt{\frac{1}{nm}}.$
Denote $u_{i}(j,k)=\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}=e_{j}e^{T}_{k}]$. Note
$\|\mathbb{E}[u_{i}X_{i}]\|$ $=\sup_{\|z\|_{2}=1,\|y\|_{2}=1}z^{T}\mathbb{E}[u_{i}X_{i}]y$
$\leq\left(\sum_{j,k}P^{2}_{jk}\right)^{1/2}\left(\sum_{j,k}z_{j}^{2}y_{k}^{2}u^{2}_{i}(j,k)\right)^{1/2}$
$\leq\max(\sigma,a)\sqrt{\frac{1}{m_{1}m_{2}}}\left(\frac{m\log^{4}d}{n}\right)^{(1+\kappa)/2}.$
Given $Mn^{\kappa}\geq m\log^{4}d$, the RHS is less than $\max(\sigma,a)\sqrt{\frac{1}{nm}}$. An alternative condition is $n\geq\log^{4/\kappa}d$. When the noise only has second moment, namely $\kappa=0$, we can derive a $\log d$-free bound given $M\geq m\log^{4}d$ and $n\geq m\log^{5}d$.
For non-symmetric noise, a $\log d$-free bound may be obtained under a better condition (i.e. $M\geq m\log^{\beta}d$ where $54\leq\beta<$ 5) by applying Theorem 13 to $\frac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])$. We do not make further efforts here since it cannot remove the condition entirely. This technical requirement will possibly be not necessary with future refinement of the sharp concentration inequalities in Brailovskaya and van Handel [2024].
Before presenting the poof of Lemma 7, we state a technical lemma.
Lemma 8.
With $\lambda$ in (4) satisfying $\lambda\geq 3\|\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\|$ , we can obtain
$\|P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})\|_{*}\leq 2\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{*}.$
Proof.
As a solution of a convex program, by Proposition 1.3 in Bubeck [2014] , $\hat{A}_{H}$ satisfies:
$\langle\nabla L(\hat{A}_{H})+\lambda Z,\bar{A}-A_{0}\rangle\leq 0.$
where $Z\in\partial\|\hat{A}_{H}\|_{*}$. Since $\|\cdot\|_{*}$ is convex, we have
$\langle Z,A_{0}-\hat{A}_{H}\rangle\leq\|A_{0}\|_{*}-\|\hat{A}_{H}\|_{*}.$
By the convexity of $L(\cdot)$,
$\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle\geq 0.$
Combining the results above, we obtain:
$0\leq\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle\leq\lambda(\|A_{0}\|_{*}-\|\hat{A}_{H}\|_{*})+\|\nabla L(A_{0})\|\|\hat{A}_{H}-A_{0}\|_{*}.$ (12)
Note
$\|\hat{A}_{H}\|_{*}$ $=\|P_{A_{0}}(\hat{A}_{H}-A_{0})+P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})+P_{A_{0}}(A_{0})\|_{*}$
$\geq\|P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})\|_{*}+\|P_{A_{0}}(A_{0})\|_{*}-\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{*}.$
Then, we get
$0\leq\frac{4}{3}\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{*}-\frac{2}{3}\|P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})\|_{*},$
which leads to the statement of the lemma. x∎
Proof Lemma 7.
As a solution of the convex program (4), by (12), $\hat{A}_{H}$ satisfies:
$\langle\nabla L(\hat{A}_{H})-L(A_{0}),\hat{A}_{H}-A_{0}\rangle\leq\frac{4}{3}\lambda\|\hat{A}_{H}-A_{0}\|_{*}.$
From Wainwright [2019] Chapter 10, we know, for a matrix $B$,
$\|P_{A_{0}}(B)\|_{*}\leq\sqrt{2\text{rank}(A_{0})}\|B\|_{F}.$
By Lemma 8,
$\|\hat{A}_{H}-A_{0}\|_{*}\leq 3\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{*}\leq 3\sqrt{2r}\|\hat{A}_{H}-A_{0}\|_{F}.$
By Lemma 9, assuming $\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}\leq s^{2}$, there is a constant $c_{3}$ such that the following holds with probability at least $1-\frac{1}{d}$,
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)\leq 4\sqrt{2r}\lambda\|\hat{A}_{H}-A_{0}\|_{F}+\dfrac{64a^{2}s^{2}}{\tau^{2}}$
$+c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{*}\mathbb{E}\|R\|+\frac{\mu a^{2}\log d}{n}\Bigg{)},$ (13)
where we use $\|\hat{A}_{H}-A_{0}\|_{\infty}\leq 2a$. On the RHS of (13), it holds
$\lambda\sqrt{r}\|\hat{A}_{H}-A_{0}\|_{F}\leq 40\mu\lambda^{2}rm_{1}m_{2}+\frac{1}{40}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}.$
By $\|\hat{A}_{H}-A_{0}\|_{\infty}\leq 2a$, we obtain
$\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{*}\mathbb{E}\|R\|$ $\leq 2a\sqrt{r}\|\hat{A}_{H}-A_{0}\|_{F}\mathbb{E}\|R\|$
$\leq 40c_{3}\mu a^{2}\mathbb{E}^{2}\|R\|rm_{1}m_{2}+\frac{1}{10c_{3}}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}.$
Then we obtain
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)\leq 160\sqrt{2}\mu\lambda^{2}rm_{1}m_{2}+\frac{1}{5}\frac{\|\hat{A}_{H}-A_{0}\|^{2}_{F}}{\mu m_{1}m_{2}}+\dfrac{64a^{2}s^{2}}{\tau^{2}}$
$+c_{3}\Bigg{(}40c_{3}\mu a^{2}\mathbb{E}^{2}\|R\|rm_{1}m_{2}+\frac{4\mu a^{2}\log d}{n}\Bigg{)},$
If $n\geq 160m\log^{4}d$, we have $\frac{4\sigma^{2}}{\tau^{2}}\leq\frac{1}{10}$ and $\frac{64a^{2}}{\tau^{2}}\leq\frac{2}{5}$, which leads to $\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\frac{4\sigma^{2}}{\tau^{2}}\right)\geq\frac{4}{5}\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}$. Then we obtain that there is a constant $c_{4}$ such that the following holds with probability at least $1-\frac{1}{d}$,
$\frac{3}{5}\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq c_{4}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{5}s^{2}.$
The above implies
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)+\frac{2}{3}s^{2}.$
Denote $\delta=\frac{5c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|R\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$. We take $s^{2}=4\delta$. If $\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}>s^{2}$, we can take $\tilde{A}_{H}=A_{0}+\eta(\hat{A}_{H}-A_{0})$ such that $\|\tilde{A}_{H}-A_{0}\|_{w(F)}=\|\eta(\hat{A}_{H}-A_{0})\|_{w(F)}=s$. Note that $\tilde{A}_{H}$ also satisfies
$\|\tilde{A}_{H}-A_{0}\|_{*}\leq 3\sqrt{r}\|\tilde{A}_{H}-A_{0}\|_{F}.$
By Lemma F.2 in Fan et al. [2018], we have
$\langle\nabla L(\tilde{A}_{H})-\nabla L(A_{0}),\tilde{A}_{H}-A_{0}\rangle$ $\leq\eta\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle$
$\leq 4\lambda\|P_{A_{0}}(\tilde{A}-A_{0})\|_{*}$
Then, we can repeat the previous steps and Lemma 9 also applies to $\tilde{A}_{H}-A_{0}$. Finally we get that
$\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\delta+\frac{2}{3}s^{2}\leq\frac{11}{12}s^{2},$
which contradicts with that $\|\tilde{A}_{H}-A_{0}\|_{w(F)}^{2}=s^{2}$. Thus, we get
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\leq\frac{20c_{4}}{3}\left(\mu(\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|)rm_{1}m_{2}+\frac{\mu\log d}{n}\right)$
and
$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}(\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|)rm_{1}m_{2}+\frac{\mu^{2}\log d}{n}.$
4.2 Proof of restricted strong convexity for Huber loss function
Lemma 9.
Assume $n\geq m\log^{4}d$ and $\|\hat{A}_{H}-A_{0}\|_{w(F)}\leq s$. With probability at least $1-\frac{1}{d}$, we have
$\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle$ $\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}\left(\frac{9}{10}-\dfrac{4\sigma^{2}}{\tau^{2}}\right)-\dfrac{16\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}s^{2}}{\tau^{2}}$
$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{*}\mathbb{E}\|R\|+\frac{\mu\|\hat{A}_{H}-A_{0}\|_{\infty}^{2}\log d}{n}\Bigg{)},$ (14)
where $c_{3}$ is a universal constant.
Proof of Lemma 9.
Consider events $H_{i}=\{|\langle X_{i},\hat{A}_{H}-A_{0}\rangle|\leq\tau/2\}\cap\{|\xi_{i}|\leq\tau/2\}$. We have the following inequality:
$\quad\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle$
$=\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\phi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle$
$\geq\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(Y_{i}-\langle X_{i},A_{0}\rangle)-\phi_{\tau}(Y_{i}-\langle X_{i},\hat{A}_{H}\rangle))\langle X_{i},\hat{A}_{H}-A_{0}\rangle\textbf{1}_{H_{i}}$
$=\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{H}-A_{0}\rangle^{2}\textbf{1}_{H_{i}}.$
Consider function $G_{R}(x):\mathbb{R}\rightarrow\mathbb{R}$,
$G_{R}(x)=\begin{cases}x^{2},|x|\leq R,\\ (x-2R\cdot\operatorname{sign}(x))^{2},R\leq|x|\leq 2R,\\ 0,\text{otherwise}.\end{cases}$
$G_{R}(x)$ has the following properties:
  • 1.
    $x^{2}\textbf{1}_{|x|\leq 2R}\geq G_{R}(x)\geq x^{2}\textbf{1}_{|x|\leq R}$.
  • 2.
    $G_{R}(x)$ is Lipschitz with $|G_{R}(x)-G_{R}(y)|\leq 2R|x-y|$ since $|G_{R}^{\prime}(x)|\leq 2R$. Moreover, if we let $|x|,|y|\leq a$, $|G_{R}(x)-G_{R}(y)|\leq 2a|x-y|$.
Assume $\|\hat{A}_{H}-A_{0}\|_{w(F)}\leq s$. Then by Property 1 of $G_{R}(\cdot)$,
$\quad\leavevmode\nobreak\ \left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle^{2}\textbf{1}_{H_{i}}$
$=\left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle^{2}\textbf{1}\left(\dfrac{\langle X_{i},\hat{A}_{H}-A_{0}\rangle}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\leq\dfrac{\tau}{2\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right)\textbf{1}\left(|\xi_{i}|\leq\tau/2\right)$
$\geq\left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle^{2}\textbf{1}\left(\dfrac{\langle X_{i},\hat{A}_{H}-A_{0}\rangle}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\leq\dfrac{\tau}{2s}\right)\textbf{1}_{\chi_{i}}$
$\geq G_{\frac{\tau}{4s}}\left(\left\langle X_{i},\frac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle\right)\textbf{1}_{\chi_{i}},$
where $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$. Hence, we have
$\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{H}-A_{0}\rangle^{2}\textbf{1}_{H_{i}}$ $=\dfrac{\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}}{n}\sum_{i=1}^{n}\left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle^{2}\textbf{1}_{H_{i}}$
$\geq\|\hat{A}_{H}-A_{0}\|^{2}_{w(F)}\cdot\left(\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}\left(\left\langle X_{i},\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right\rangle\right)\textbf{1}_{\chi_{i}}\right).$
For a matrix $A\in\mathbb{R}^{m_{1}\times m_{2}}$, we define
$f(A):=\mathbb{E}_{X_{i},\xi_{i}}[G_{\frac{\tau}{4s}}(X_{i},A)\textbf{1}_{\chi_{i}}].$
We have,
$f(A)$ $\geq\mathbb{E}[\langle X_{i},A\rangle^{2}]\textbf{1}\{|\langle X_{i},A\rangle|\leq\tau/4s\}\textbf{1}_{\chi_{i}}$
$\geq\mathbb{E}\langle X_{i},A\rangle^{2}-\mathbb{E}\langle X_{i},A\rangle^{2}\textbf{1}\{|\langle X_{i},A\rangle|>\tau/4s\}-\mathbb{E}\langle X_{i},A\rangle^{2}\textbf{1}\{|\xi_{i}|>\tau/2\}$
$\geq\|A\|_{w(F)}^{2}-\dfrac{16s^{2}\|A\|^{2}_{\infty}}{\tau^{2}}\|A\|_{w(F)}^{2}-\dfrac{4\sigma^{2}}{\tau^{2}}\|A\|_{w(F)}^{2}$
$=\|A\|_{w(F)}^{2}\left(1-\dfrac{16s^{2}\|A\|_{\infty}^{2}}{\tau^{2}}-\dfrac{4\sigma^{2}}{\tau^{2}}\right).$
Thus
$\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}f\left(\dfrac{\hat{A}_{H}-A_{0}}{\|\hat{A}_{H}-A_{0}\|_{w(F)}}\right)\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}(1-\dfrac{4\sigma^{2}}{\tau^{2}})-\dfrac{16\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}s^{2}}{\tau^{2}}.$ (15)
Now we introduce a useful lemma.
Lemma 10.
Assume $\|A\|_{w(F)}^{2}=1$. Then there are universal constants $c_{0},c_{1},c_{2}$ such that
$\left|\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)\textbf{1}_{\chi_{i}}-f(A)\right|\leq 8c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$ (16)
for all such $A$, uniformly with probability at least $1-\frac{1}{d}$. Here $\mathcal{R}=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}$ and $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
By (15) and (16), we get
$\quad\langle\nabla L(\hat{A}_{H})-\nabla L(A_{0}),\hat{A}_{H}-A_{0}\rangle\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}(1-\dfrac{4\sigma^{2}}{\tau^{2}})-\dfrac{16a^{2}s^{2}}{\tau^{2}}$
$-\Bigg{(}8c_{0}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|\hat{A}_{H}-A_{0}\|_{\infty}\|\hat{A}_{H}-A_{0}\|_{w(F)}\sqrt{\frac{\mu\log d}{n}}$
$+8c_{2}\frac{\mu\|\hat{A}_{H}-A_{0}\|_{\infty}^{2}\log d}{n}\Bigg{)}$
$\geq\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}(1-\dfrac{4\sigma^{2}}{\tau^{2}})-\dfrac{16\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}s^{2}}{\tau^{2}}$
$-c_{3}\Bigg{(}\|\hat{A}_{H}-A_{0}\|_{*}\|\hat{A}_{H}-A_{0}\|_{\infty}+\frac{\mu\|\hat{A}_{H}-A_{0}\|^{2}_{\infty}\log d}{n}\Bigg{)}-\frac{\|\hat{A}_{H}-A_{0}\|_{w(F)}^{2}}{10}$ (17)
where in the last step, we apply the basic inequality ($2ab\leq\frac{a^{2}}{c^{2}}+c^{2}b^{2}$) and choose a suitable constant $c_{3}$. By (17), we get (14).
Proof of Lemma 10.
At first, we introduce a classical concentration inequality for empirical processes, which can be found in many textbooks such as Wainwright [2019].
Lemma 11 (Theorem 3.27 in Wainwright [2019]).
Let $Z_{i},1\leq i\leq n$, be independent but not necessarily identically distributed random variables taking values in measurable spaces $\mathcal{Z},1\leq i\leq n$. $\mathscr{F}$ is a function class. Consider the random variable
$Z=\sup_{f\in\mathscr{F}}\frac{1}{n}\sum_{i=1}^{n}f(Z_{i}).$
If $\|f\|_{\infty}\leq b$ for all $f\in\mathscr{F}$, we have the following inequality, for $t>0$,
$P(Z\geq c_{0}E[Z]+c_{1}\tilde{\sigma}\sqrt{t}+c_{2}bt)\leq\exp\left(-nt\right)$ (18)
where $\tilde{\sigma}^{2}=\sup_{f\in\mathscr{F}}\frac{1}{n}\sum_{i=1}^{n}Ef^{2}(Z_{i})$ and $c_{0},c_{1},c_{2}$ are some universal constants.
For fixed $(\alpha,\rho)\in\mathbb{R}^{2}$, we define the space:
$S(\alpha,\rho)=\{A:\|A\|_{w(F)}^{2}=1,\|A\|_{*}\leq\rho,\|A\|_{\infty}\leq\alpha\}.$
Denote
$F_{A}(X_{i},\xi_{i})=G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)\textbf{1}_{\chi_{i}}.$
for $A\in S(\alpha,\rho)$. Recall $f(A)=\mathbb{E}_{X_{i},\xi_{i}}[F_{A}(X_{i},\xi_{i})]$. Since $x^{2}\geq x^{2}\textbf{1}_{|x|\leq 2R}\geq G_{R}(x)$, for $A\in S(\alpha,\rho)$,
$\quad\leavevmode\nobreak\ \mathbb{E}[F^{2}_{A}(X_{i},\xi_{i})]-f^{2}(A)$ $\leq\mathbb{E}[F^{2}_{A}(X_{i},\xi_{i})]$
$\leq\mathbb{E}[G^{2}_{\tau/4s}(\langle X_{i},A\rangle)]$
$\leq\alpha^{2}\|A\|_{w(F)}^{2}$
$\leq\alpha^{2},$
which implies $\tilde{\sigma}\leq\alpha$ in (18). Consider
$Z(\alpha,\rho)=\sup_{A\in S(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|$
We have
$\mathbb{E}[Z(\alpha,\rho)]$ $\stackrel{{\scriptstyle\text{i}}}{{\leq}}\mathbb{E}\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}F_{A}(X_{i},\xi_{i})\right|$
$\stackrel{{\scriptstyle\text{ii}}}{{\leq}}2\alpha\mathbb{E}\sup_{A\in\Omega(\alpha,\rho)}\left|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\langle X_{i}\textbf{1}_{\chi_{i}},A\rangle\right|$
$\stackrel{{\scriptstyle\text{iii}}}{{\leq}}4\alpha\rho\mathbb{E}\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|.$
In step (i) we use symmetrization technique. By Property (ii) of $G_{R}(\cdot)$ and $\|A\|_{\infty}\leq\alpha$, $G_{\frac{\tau}{4s}}(\cdot)$ is Lipschitz with Lipschitz constant less than $2\alpha$. Then, in step (ii) we apply Talagrand-Ledoux contraction inequality. Step (iii) is due to $|\langle A,B\rangle|\leq\|A\|\|B\|_{*}$.
We take $t=\frac{2\mu\log d}{n}$ in (18). Since $\tilde{\sigma}\leq\alpha$ and $\|F_{A}(X_{i},\xi_{i})\|_{\infty}\leq\alpha^{2}$, with probability at least $1-\exp(-2\mu\log d)$,
$Z(\alpha,\rho)\leq 4c_{0}\alpha\rho\mathbb{E}\|\mathcal{R}\|+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\alpha^{2}\log d}{n}.$ (19)
Note that under Assumption 3,
$\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}\|A\|_{w(F)}=\sqrt{\mu m_{1}m_{2}}.$
and
$\|A\|_{F}\geq\|A\|_{w(F)}=1.$
Then, we have
$\|A\|_{*}\leq\sqrt{m}\|A\|_{F}\leq\sqrt{\mu d^{3}}$
and
$1\leq\|A\|_{F}\leq\|A\|_{*}.$
Without loss of generality, we can assume $\|A\|_{*}\in[1,\sqrt{\mu d^{3}}]$. Similarly, we have
$\|A\|_{\infty}\leq\|A\|_{F}\leq\sqrt{\mu m_{1}m_{2}}$
and
$\|A\|_{\infty}\geq\frac{\|A\|_{F}}{\sqrt{m_{1}m_{2}}}\geq\frac{1}{d}.$
Consider
$\tilde{S}(k,l)=\{A;\|A\|_{w(F)}=1,2^{k-1}\leq\|A\|_{\infty}\leq 2^{k},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ 2^{l-1}\leq\|A\|_{*}\leq 2^{l}\}$
Every matrix $A$ with $\|A\|_{w(F)}=1$ belongs to some $\tilde{S}(k,l)$ with $k=\lfloor\log_{2}\frac{1}{d}\rfloor,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d^{2})\rceil$ and $l=1,\ldots,\lceil\log_{2}\frac{1}{2}(\mu d^{3})\rceil$. Thus the total number of the pairs $(k,l)$ is less than $\Lambda^{2}=9\log_{2}^{2}(\mu d)$.
Denote the events $\mathcal{E}$ that a matrix $A$ violates the following relation
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 16c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$ (20)
and denote $\mathcal{E}_{k,l}$ the event that there exists a matrix in $\tilde{S}(k,l)$ violating the following relation
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 4c_{0}2^{k}2^{l}\mathbb{E}\|\mathcal{R}\|+\sqrt{2}c_{1}2^{k}\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}.$
$\mathcal{E}$ is contained in the union $\cup_{k,l}\mathcal{E}_{k,l}$. Indeed, letting $\alpha=2^{k}$ and $\rho=2^{l}$, for a matrix $A$ in $\tilde{S}(k,l)$, we have
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|$ $\geq 16c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$
$\geq 16c_{0}2^{k-1}2^{l-1}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}2^{k-1}\sqrt{\frac{\mu\log d}{n}}+16c_{2}2^{2(k-1)}\frac{\mu\log d}{n}$
$=4c_{0}\alpha\rho+\sqrt{2}\alpha\sqrt{\frac{\mu\log d}{n}}+4c_{2}\alpha^{2}\frac{\mu\log d}{n}.$
Thus, by (19),
$\mathbb{P}(\mathcal{E})\leq\mathbb{P}(\cup_{k,l}\mathcal{E}_{k,l})\leq\Lambda^{2}\exp(-2\mu\log d)\leq\frac{9\log_{2}^{2}\mu d}{d^{\mu}}\cdot\frac{1}{d^{\mu}}\leq\frac{1}{d}$
for sufficiently large $d$. ∎
4.3 Proof of Theorem 3
From Appendix A in Klopp [2014], we know the following two inequalities. Given $\lambda\geq 3\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$, we have
$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{0}\rangle^{2}\leq\frac{5}{3}\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}.$ (21)
We also have
$\|\hat{A}-A_{0}\|_{*}\leq\sqrt{72r}\|\hat{A}-A_{0}\|_{F}.$ (22)
Regarding the difference between $\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{0}\rangle^{2}$ and $\|\hat{A}-A_{0}\|^{2}_{w(F)}$, we have the following lemma.
Lemma 12.
Assume $\|A\|_{w(F)}^{2}=1$. Then there is a constant $C$ dependent on $c_{0},c_{1},c_{2}$ such that
$\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|^{2}_{w(F)}\right|\leq C\left(\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+\frac{\mu\|A\|_{\infty}^{2}\log d}{n}\right)$ (23)
for all such $A$, uniformly with probability at least $1-\frac{1}{d}$. Here $\mathcal{R}=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$.
Proof.
Denote
$Z_{\tau}(A)=\left|\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)\textbf{1}_{\chi_{i}}-f(A)\right|,$
$Z=\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|^{2}_{w(F)}\right|,$
$\kappa(A)=8c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+8c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}.$
Then, we consider the variables
$Y_{\tau}=\sup_{\|A\|_{w(F)}=1}\frac{Z_{\tau}(A)}{\kappa(A)},$
$Y=\sup_{\|A\|_{w(F)}=1}\frac{Z}{\kappa(A)}.$
Given $\|A\|_{w(F)}=1$, we have that $\kappa(A)$ is uniformly lower bounded away from zero (see the proof of Lemma 10). And we note
$\left|\dfrac{1}{n}\sum_{i=1}^{n}G_{\frac{\tau}{4s}}(\langle X_{i},A\rangle)\textbf{1}_{\chi_{i}}-\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}\right|$ $\leq C_{1}sa^{2}\frac{1}{n}\sum_{i=1}^{n}\textbf{1}\left(|\xi_{i}|>\frac{\tau}{2}\right),$
$\left|f(A)-\|A\|^{2}_{w(F)}\right|$ $\leq C_{2}sa^{2}/\tau,$
where $C_{1},C_{2}$ are constants that are not dependent on $\tau$. Thus, we have that $Y_{\tau}$ converges to $Y$ in probability and thus in distribution. There exist constants y (dense in $\mathbb{R}$) such that
$\lim_{\tau\rightarrow\infty}\mathbb{P}(Y_{\tau}\leq y)=\mathbb{P}(Y\leq y).$
For $y\geq 1$, by Lemma 10, it holds that
$\mathbb{P}(Y_{\tau}\leq y)\geq 1-\frac{1}{d}$
for all $\tau$. There exists some $y\geq 1$ such that
$P(Y\leq y)\geq 1-\frac{1}{d},$
which implies Lemma 12. ∎
Combining Lemma 12 and (21), we obtain
$\|\hat{A}-A_{0}\|^{2}_{w(F)}$ $\leq 8c_{0}\|\hat{A}-A_{0}\|_{\infty}\|\hat{A}-A_{0}\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|\hat{A}-A_{0}\|_{w(F)}\|\hat{A}-A_{0}\|_{\infty}\sqrt{\frac{\mu\log d}{n}}$
$+8c_{2}\frac{\mu\|\hat{A}-A_{0}\|_{\infty}\log d}{n}+\frac{5}{3}\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}.$
By $\|\hat{A}-A_{0}\|_{\infty}\leq 2a$ and (22), we obtain the following inequalities:
$\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}\leq 5u\lambda^{2}rm_{1}m_{2}+\frac{1}{5}\frac{\|\hat{A}-A_{0}\|^{2}_{F}}{\mu}\leq 5u\lambda^{2}rm_{1}m_{2}+\frac{1}{5}\|\hat{A}-A_{0}\|^{2}_{w(F)},$
$\frac{\mu\|\hat{A}-A_{0}\|_{\infty}\log d}{n}\leq\frac{2\mu a\log d}{n},$
$\|\hat{A}-A_{0}\|_{\infty}\|\hat{A}-A_{0}\|_{*}\mathbb{E}\|\mathcal{R}\|\leq 5\|\hat{A}-A_{0}\|_{\infty}\sqrt{r}\|\hat{A}-A_{0}\|_{F}\mathbb{E}\|\mathcal{R}\|\leq 5\times 4\mu a^{2}rm_{1}m_{2}\mathbb{E}^{2}\|\mathcal{R}\|+\frac{1}{5}\|\hat{A}-A_{0}\|^{2}_{w(F)},$
$\|\hat{A}-A_{0}\|_{w(F)}\|\hat{A}-A_{0}\|_{\infty}\sqrt{\frac{\mu\log d}{n}}\leq\frac{1}{5}\|\hat{A}-A_{0}\|^{2}_{w(F)}+5\times 4\frac{\mu a^{2}\log d}{n}.$
Putting together the pieces, we get
$\|\hat{A}-A_{0}\|^{2}_{w(F)}\lesssim\mu m_{1}m_{2}\left(r\lambda^{2}+a^{2}\mathbb{E}^{2}\|\Sigma_{R}\|\right)+\frac{\mu a^{2}\log d}{n}.$ (24)
By (29), we obtain
$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\lesssim\sigma\sqrt{\frac{1}{nm}},$
given $n\geq m\log^{4}d\cdot(\log d+\log n)$. Then, we can take $\lambda\sim\sigma\sqrt{\frac{1}{nm}}$. By (31),
$\mathbb{E}\left\|\sum\epsilon_{i}X_{i}\right\|\lesssim\sqrt{\frac{1}{nm}}.$
Plugging the above two estimates into (24), we obtain
$\frac{\|\hat{A}-A_{0}\|^{2}_{F}}{m_{1}m_{2}}\lesssim\frac{\mu^{2}\max(a^{2},\sigma^{2})rM}{n}+\frac{\mu^{2}a^{2}\log d}{n},$
which leads to the statement of Theorem 3.
4.4 Proof of Theorem 5
In this section, we denote $\xi=(\xi_{1}/\sqrt{n},\ldots,\xi_{n}/\sqrt{n})$. From Appendix D in Klopp [2014], given
$\lambda\geq\frac{3\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|}{\|\xi\|_{2}},$
we have that
$\|\hat{A}_{S}-A_{0}\|_{*}\leq\sqrt{9r}\|\hat{A}_{S}-A_{0}\|_{F}.$ (25)
Again, from Appendix D in Klopp [2014], we have
$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}_{S}-A_{0}\rangle^{2}\leq\frac{14}{3}\lambda Q(A_{0})\sqrt{2r}\|\hat{A}_{S}-A_{0}\|_{F}+\frac{1}{4}\|\hat{A}_{S}-A_{0}\|^{2}_{w(F)},$
under condition $44\mu m_{1}m_{2}\lambda^{2}r\leq 1/$ 4. By Lemma 12, we get
$\|\hat{A}_{S}-A_{0}\|_{w(F)}^{2}$ $\leq 8c_{0}\|\hat{A}_{S}-A_{0}\|_{\infty}\|\hat{A}_{S}-A_{0}\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|\hat{A}_{S}-A_{0}\|_{w(F)}\|\hat{A}_{S}-A_{0}\|_{\infty}\sqrt{\frac{\mu\log d}{n}}$
$+8c_{2}\frac{\mu\|\hat{A}_{S}-A_{0}\|_{\infty}\log d}{n}+\frac{14}{3}\lambda\|\xi\|_{2}\sqrt{2r}\|\hat{A}_{S}-A_{0}\|_{F}+\frac{1}{4}\|\hat{A}_{S}-A_{0}\|^{2}_{w(F)}.$
Applying $\|\hat{A}_{S}-A_{0}\|_{\infty}\leq 2a$ and (25), repeat the same precesses as in Proof of Theorem 3 and we obtain
$\|\hat{A}_{S}-A_{0}\|_{w(F)}^{2}\lesssim\mu m_{1}m_{2}r\left(\|\xi\|^{2}_{2}\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|\right)+\frac{\mu\log d}{n},$ (26)
where $\mathcal{R}$ is defined as in Lemma 12. By Theorem 3.1.1 in Vershynin [2018], denoting $\xi=(\xi_{1}/n,\ldots,\xi_{n}/n)$, we get
$\mathbb{P}\left(\left|\left\|\xi\right\|_{2}-\sigma\right|\geq L_{1}\sigma t\right)\leq 2\exp(-Cnt^{2}),$
under Assumption 2. With $t=0.1/L_{1}$, we obtain that with probability at least $1-\exp(-Cn)$ ($C$ is dependent on $L_{1}$), the following holds
$0.9\sigma\leq\|\xi\|_{2}\leq 1.1\sigma.$
By (29),
$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\lesssim\sigma\sqrt{\frac{1}{nm}},$
given $n\geq m\log^{4}d\cdot(\log d+\log n)$. Thus, we can take $\lambda\sim\sqrt{\frac{1}{nm}}$. By (31),
$\mathbb{E}\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\lesssim\sqrt{\frac{1}{nm}}.$
Plugging $\lambda$ and the upper bounds for $\mathbb{E}\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|$ into (26). we obtain the convergence rate
$\frac{\|\hat{A}_{S}-A_{0}\|_{2}^{2}}{m_{1}m_{2}}\lesssim\frac{\mu^{2}\max(a^{2},\sigma^{2})rM}{n}+\frac{\mu^{2}a^{2}\log d}{n},$
where the constants are dependent on $L_{1}\sim L_{3}$. This leads to the statement of Theorem 5 as the second term above on RHS is less that the first term.
4.5 Proof of sharp bounds for spectral norms
In this section, we present advanced inequalities introduced in Brailovskaya and van Handel [2024] and how to use them. Let $Q=\sum_{i=1}^{n}Q_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ be a random matrix where $Q_{i}$ are independent random matrices with mean zero. We define
$\gamma(Q)$ $=\left(\max(\|\mathbb{E}[QQ^{T}]\|,\|\mathbb{E}[Q^{T}Q]\|)\right)^{1/2},$
$\gamma_{*}(Q)$ $=\sup_{\|y\|_{2}=\|z\|_{2}=1}\left(\mathbb{E}(y^{T}Qz)^{2}\right)^{1/2},$
$g(Q)$ $=\|\operatorname{Cov}(Q)\|^{1/2},$
$R(Q)$ $=\left\|\max_{1\leq i\leq n}\|Q_{i}\|\right\|_{\infty},$
where $\operatorname{Cov}(Q)$ is the covariance matrix of all the entries of $Q$ and its dimension is $m_{1}m_{2}\times m_{1}m_{2}$.
Theorem 13 (Corollary 2.17 in Brailovskaya and van Handel [2024]).
With the above notations, there exist a universal constant $C$ such that
$\mathbb{P}(\|Q\|\geq 2\gamma(Q)+C(g(Q)^{1/2}\gamma(Q)^{1/2}\log^{3/4}d+\gamma_{*}(Q)t^{1/2}+R(Q)^{1/3}\gamma(Q)^{2/3}t^{2/3}+R(Q)t))\leq de^{-t}$ (27)
and a universal constant $C$ such that
$E\|Q\|\leq 2\gamma(Q)+C(g(Q)^{1/2}\gamma(Q)^{1/2}\log^{3/4}d+R(Q)^{1/3}\gamma(Q)^{2/3}\log^{2/3}d+R(Q)\log d).$ (28)
Remark.
The matrix inequalities in Brailovskaya and van Handel [2024] are formulated for $Q_{i}\in\mathbb{R}^{d\times d}$ (not necessarily self-adjoint), and can be extended to rectangular matrices in the above form. See Remark 2.1 in Brailovskaya and van Handel [2024].
Theorem 13 requires $Q_{i}$ to have bounded spectral norms, which is too restrictive in the context of matrix completion. To leverage Theorem 13, we need to truncate the noise variables. Brailovskaya and van Handel [2024] also provides concentration inequalities for the unbounded case (Theorem 2.8) that is derived by truncation techniques. However, applying Theorem 2.8 in Brailovskaya and van Handel [2024] could result in worse sample size conditions as Brailovskaya and van Handel [2024] considers general cases. For matrix completion, we need derive specially tailored concentration inequalities from Theorem 13 (essentially from Theorem 2.6 in Brailovskaya and van Handel [2024]).
Lemma 14.
(1) Assume the sampling matrices $X_{i}$ satisfy Assumption 3 and $\xi_{i}$ are $\psi_{\alpha}$ random variables with $\mathbb{E}[\xi_{i}]=0$, $\mathbb{E}[\xi_{i}^{2}|X_{i}]\leq\sigma^{2}$ and uniformly bounded $\psi_{\alpha}$ norms
$\left\|[\xi_{i}|X_{i}]\right\|_{\psi_{\alpha}}\leq L_{1}\sigma.$
The following inequality holds with probability at least $1-\frac{1}{d}$,
$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C\sigma\sqrt{\dfrac{1}{nm}},$ (29)
given $n\geq m\log^{4}d\cdot(\log^{2/\alpha}n+\log^{2/\alpha}d)$. In particular, for sub-Gaussian noise, (29) holds given $n\geq m\log^{4}d\cdot(\log n+\log d)$. Here $C$ is a constant dependent on $L_{1}\sim L_{3}$.
(2)Assume $\xi_{i}$ satisfy Assumption 1 and are symmetric and $X_{i}$ satisfy Assumption 3. Taking $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$, the following inequality holds with probability at least $1-\frac{1}{d}$,
$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\right\|\leq C\max(\sigma,a)\sqrt{\dfrac{1}{nm}}.$ (30)
where $C$ is dependent on $L_{1}$ and $L_{3}$.
(3) If $n\geq m\log^{4}d$, we have the following result:
$\mathbb{E}\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C\sqrt{\frac{1}{nm}},$ (31)
where $C$ is dependent on $L_{2}$ and $L_{3}$.
(4) If $n\geq m\log^{4}d$, we have the following inequality:
$\mathbb{E}\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|\leq C\sqrt{\frac{1}{nm}},$ (32)
where $\xi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
Proof of Lemma 14.
Since Theorem 13 requires $\|Q_{i}\|$ to be bounded, we need to truncate $\xi_{i}$ to $\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)$ for some $\tau\geq 0$. Define $\bar{\xi}_{i}=\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)-\mathbb{E}[\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)|X_{i}]$. Note $\mathbb{E}[\bar{\xi}_{i}]=0$ and $\mathbb{E}[\bar{\xi}_{i}^{2}]\leq\sigma^{2}$. We now compute the four parameters for $Q=\frac{1}{n}\sum_{i=1}^{n}\bar{\xi}_{i}X_{i}$. Define the event
$\mathcal{E}_{i}(j,k)=\{X_{i}=e_{j}(m_{1})e_{k}(m_{2})^{T}\}.$
1. $\gamma(Q)\leq\sqrt{\frac{L_{2}\sigma^{2}}{nm}}$. Note $\mathbb{E}[\bar{\xi}^{2}_{i}X_{i}X_{i}^{T}]$ is a diagonal matrix. The elements on the diagonal line are $\sum_{k=1}^{m_{2}}P_{jk}\mathbb{E}[\bar{\xi}_{i}^{2}|\mathcal{E}_{i}(j,k)],j=1,2,\ldots,m_{1}$, which are less than $\frac{L_{2}\sigma^{2}}{m}$. The same bound holds for $\mathbb{E}[\bar{\xi}_{i}^{2}X_{i}^{T}X_{i}]$. Since
$\mathbb{E}[QQ^{T}]=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}[\bar{\xi}_{i}^{2}X_{i}X_{i}^{T}],$
we obtain $\|E[QQ^{T}]\|\leq\frac{L_{2}\sigma^{2}}{nm}$. Similarly, $\|\mathbb{E}[Q^{T}Q]\|\leq\frac{L_{2}\sigma^{2}}{nm}$. Thus $\gamma(Q)\leq\sqrt{\frac{L_{2}\sigma^{2}}{nm}}$.
2. $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$. For $y\in\mathbb{R}^{m_{1}},z\in\mathbb{R}^{m_{2}}$ with $\|y\|_{2}=\|z\|_{2}=1$, we have the following result:
$\mathbb{E}(y^{T}\bar{\xi}_{i}X_{i}z)^{2}$ $=\sum_{j,k}P_{jk}\mathbb{E}[\bar{\xi}_{i}^{2}|\mathcal{E}_{i}(j,k)]y^{2}_{j}z^{2}_{k}$
$\leq\sigma^{2}\sum_{j,k}P_{jk}y^{2}_{j}z^{2}_{k}$
$\leq\dfrac{L_{3}\sigma^{2}}{m\log^{3}d}.$
where we use $\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}y^{2}_{j}z^{2}_{k}=\|y\|_{2}^{2}\cdot\|z\|_{2}^{2}=1$ and $\max P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$. Since $\bar{\xi}_{i}X_{i}$ are independent and mean-zero, $\mathbb{E}(y^{T}Qz)^{2}=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}\left(y^{T}\bar{\xi}_{i}X_{i}z\right)^{2}\leq\frac{L_{1}L_{3}\sigma^{2}}{nm\log^{3}d}$. Thus $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$.
3. $g(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$. Since $\bar{\xi}_{i}X_{i}$’s are independent, $\operatorname{Cov}(Q)=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\operatorname{Cov}(\bar{\xi}_{i}X_{i})$. Define $\textbf{1}_{(j,k)}=\textbf{1}(X_{i}=e_{j}(m_{1})e_{k}(m_{2})^{T})$ (for simplicity of notation, we do not include a subscript i) and denote $I(j,k)=\bar{\xi}_{i}\textbf{1}_{(j,k)}$. Note the following facts
$\mathbb{E}[I(j,k)]=P_{jk}\mathbb{E}[\bar{\xi}_{i}|\mathcal{E}_{i}(j,k)]=0,$
$I(j,k)I(l,h)=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).$
Thus $\text{Cov}\big{(}I(j,k),I(l,h)\big{)}=\mathbb{E}[I(j,k)I(l,h)]-\mathbb{E}[I(j,k)]E[I(l,h)]=0$ for $(j,k)\neq(l,h)$. Then we obtain
$\text{Cov}(\bar{\xi}_{i}X_{i})\Big{(}(j,k),(l,h)\Big{)}=\begin{cases}\mathbb{E}[\bar{\xi}^{2}_{i}|\mathcal{E}_{i}(j,k)]P_{jk}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)=(l,h),\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).\end{cases}$
Since $P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$, $\|\operatorname{Cov}(\bar{\xi}_{i}X_{1})\|\leq\frac{L_{3}\sigma^{2}}{m\log^{3}d}$. Thus, $g(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$ holds.
4.$R(Q)\leq\frac{2\tau}{n}$. Obviously, $\|\bar{\xi}_{i}X_{i}\|=|\bar{\xi}_{i}|$, and $\max_{1\leq i\leq n}|\bar{\xi}_{i}|\leq 2\tau$.
Now, we can plug the four estimates in (27). Taking $t=3\log d$, we get that with probability at least $1-\frac{1}{d^{2}}$, the following holds
$\|Q\|$ $\leq 2\sqrt{\frac{\sigma^{2}}{nm}}$
$+C\Bigg{(}\left(\frac{\sigma^{2}}{nm}\right)^{1/2}\cdot\left(\dfrac{\log^{3/4}d}{\log^{3/4}d}\right)+\left(\frac{\sigma^{2}}{nm}\right)^{1/2}\cdot\left(\dfrac{\log^{1/2}d}{\log^{3/2}d}\right)$
$+\left(\frac{\log^{2}d\cdot\tau}{n}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}+\dfrac{\log d\cdot\tau}{n}\Bigg{)},$ (33)
where $C$ is dependent on $L_{2}\sim L_{3}$. We can see the quantities in the first line and the second line of RHS of (33) are less than $C\sqrt{\frac{\sigma^{2}}{nm}}$. When $\xi_{i}$ are $\psi_{\alpha}$ random variables, we take $\tau=C_{\alpha}\sigma(\log^{1/\alpha}n+\log^{1/\alpha}d)$. By (35), with $C_{\alpha}$ properly chosen,
$\mathbb{P}\left(\max_{1\leq i\leq n}|\xi_{i}|\geq C_{\alpha}\sigma\left(\log^{1/\alpha}n+\log^{1/\alpha}d\right)\right)\leq\frac{1}{d^{2}}.$
To make the quantities in the last line of (33) be less than $C\sqrt{\frac{\sigma^{2}}{nm}}$, we need
$\left(\frac{\sigma\log^{2}d\cdot(\log^{1/\alpha}n+\log^{1/\alpha}d)}{n}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}\leq\sqrt{\dfrac{\sigma^{2}}{nm}},$
$\dfrac{\sigma\log d\cdot(\log^{1/\alpha}n+\log^{1/\alpha}d)}{n}\leq\sqrt{\dfrac{\sigma^{2}}{nm}}.$
The above conditions imply
$n$ $\geq m\log^{4}d\cdot(\log^{2/\alpha}n+\log^{2/\alpha}d),$
$n$ $\geq m\log^{2}d\cdot(\log^{2/\alpha}n+\log^{2/\alpha}d).$
On $\{\max_{1\leq i\leq n}|\xi_{i}|\leq\tau\}$,
$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|$ $=\left\|Q+\dfrac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\xi_{i}\textbf{1}(|\xi_{i}|\leq\tau)]X_{i}\right\|$
$\leq\|Q\|+\left\|\dfrac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\xi_{i}\textbf{1}(|\xi_{i}|>\tau)]X_{i}\right\|$
$\leq\|Q\|+\frac{1}{n}\sum_{i=1}^{n}|\mathbb{E}[\xi_{i}\textbf{1}(|\xi_{i}|>\tau)|X_{i}]|,$
where we use $\mathbb{E}[\xi_{i}\textbf{1}(\xi_{i}\leq\tau)]=\mathbb{E}[\xi_{i}\textbf{1}(\xi_{i}>\tau)]$ since $\mathbb{E}[\xi_{i}]=0$. For $\tau=C_{\alpha}\sigma(\log^{1/\alpha}n+\log^{1/\alpha}d)$ and a random variable $Y$ with $\|Y\|_{\psi_{\alpha}}\leq L_{1}\sigma$ , we can find suitable large constant $C_{\alpha}$ dependent on $L_{1}$ such that the following holds
$\mathbb{E}[Y\textbf{1}(|Y|>\tau)]$ $\leq\mathbb{E}[|Y|\textbf{1}(|Y|>\tau)]$
$\leq\mathbb{E}^{1/2}[Y^{2}]P^{1/2}(|Y|>\tau)$
$\leq\sigma\exp\left(-2\max(\log^{1/\alpha}n,\log^{1/\alpha}d)^{\alpha}\right)$
$\leq\frac{\sigma}{n}\leq\sigma\sqrt{\frac{1}{nm}},$
given $n\geq m$. Combining the results, we obtain
$\mathbb{P}\left(\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\geq C_{\alpha}\sigma\sqrt{\frac{1}{nm}}\right)$ $\leq\mathbb{P}\left(\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\geq C_{\alpha}\sigma\sqrt{\frac{1}{nm}},\max_{1\leq i\leq n}|\xi_{i}|\leq\tau\right)+\mathbb{P}\left(\max_{1\leq i\leq n}|\xi_{i}|>\tau\right)$
$\leq\frac{1}{d},$
given $n\geq m\log^{4}d\cdot(\log^{2/\alpha}n+\log^{2/\alpha}d)$. In particular, for sub-Gaussian noise, we have that
$\mathbb{P}\left(\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C\sigma\sqrt{\frac{1}{nm}}\right)\geq 1-\frac{1}{d}$
given $n\geq\log^{4}d\cdot(\log n+\log d)$. So far, we conclude (29)
To prove the dimension free bound for $\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}$, we denote $Q_{i}=\frac{1}{n}\phi_{\tau}(\xi_{i})X_{i}$ and $Q=\sum_{i=1}^{n}Q_{i}$. Since $\xi_{i}$ are symmetric, $\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}]=0$. By (37), we obtain $\mathbb{E}[\psi(\xi_{i})^{2}|X_{i}]\leq\sigma^{2}$, $|\phi_{\tau}(\xi_{i})|\leq\tau$. Therefore, the estimates we have obtained earlier for the four parameters still apply. By (28), with $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$, we need to impose conditions to $n$ such that in (33)
$\left(\frac{\log^{2}d}{n}\cdot\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}\right)^{1/3}\cdot\left(\frac{\sigma^{2}}{nm}\right)^{1/3}\leq\max(\sigma,a)\sqrt{\dfrac{1}{nm}},$
$\dfrac{\log d}{n}\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}\leq\max(\sigma,a)\sqrt{\dfrac{1}{nm}}.$
Given $n\geq m\log^{4}d$, the upper bound of $\|\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\|$ is free of the $\log d$ factor. We have proved (30).
Similarly, let $Q=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$. Since $\epsilon_{i}\leq 1$, we have $R(Q)=\frac{1}{n}$. We have already estimated the four parameters above. By (28), it remains to ensure that
$\left(\frac{\log^{2}d}{n}\right)^{1/3}\cdot\left(\frac{1}{nm}\right)^{1/3}\leq C\sqrt{\dfrac{1}{nm}},$
$\dfrac{\log d}{n}\leq C\sqrt{\dfrac{1}{nm}},$
which leads to the condition $n\geq m\log^{4}d.$ This completes the proof of (31).
Hereafter, we use the notation $Q_{i}=\frac{1}{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}$ and $Q=\sum_{i=1}^{n}Q_{i}$.
1. $\gamma(Q)\leq\sqrt{\frac{L_{2}}{nm}}$. Note $Q_{i}Q_{i}^{T}=\frac{1}{n^{2}}\textbf{1}_{\chi_{i}}X_{i}X_{i}^{T}$. $\mathbb{E}[\frac{1}{n^{2}}\textbf{1}_{\chi_{i}}X_{i}X_{i}^{T}]$ is a diagonal matrix with diagonal elements $\sum_{k=1}^{m_{2}}\mathbb{P}(\chi_{i}|\mathcal{E}_{i}(j,k))P_{jk}$. Thus, we have $\|Q_{i}Q_{i}^{T}\|\leq\frac{L_{2}}{n^{2}m}$ and $\|QQ^{T}\|\leq\frac{L_{2}}{nm}$. Similarly, we can obtain $\|QQ^{T}\|\leq\frac{L_{2}}{nm}$ and therefore $\gamma(Q)\leq\sqrt{\frac{L_{2}}{nm}}$.
2. $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}}{nm\log^{3}d}}$. For $y\in\mathbb{R}^{m_{1}},z\in\mathbb{R}^{m_{2}}$ with $\|y\|_{2}=\|z\|_{2}=1$, we have the following result:
$\mathbb{E}(y^{T}\bar{\xi}_{i}X_{i}z)^{2}$ $=\sum_{j,k}P_{jk}\mathbb{P}(\textbf{1}_{\chi_{i}}|\mathcal{E}_{i}(j,k))y^{2}_{j}z^{2}_{k}$
$\leq\sum_{j,k}P_{jk}y^{2}_{j}z^{2}_{k}$
$\leq\dfrac{L_{3}}{m\log^{3}d}.$
where we use $\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}y^{2}_{j}z^{2}_{k}=\|y\|_{2}^{2}\cdot\|z\|_{2}^{2}=1$ and $\max P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$. Since $Q_{i}$ are independent and mean-zero, $\mathbb{E}(y^{T}Qz)^{2}=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{E}\left(y^{T}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}z\right)^{2}\leq\frac{L_{3}}{nm\log^{3}d}$. Thus $\gamma_{*}(Q)\leq\sqrt{\frac{L_{3}\sigma^{2}}{nm\log^{3}d}}$.
3. $g(Q)\leq\sqrt{\frac{L_{3}}{nm\log^{3}d}}$. Since $Q_{i}$ are independent, $\operatorname{Cov}(Q)=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\operatorname{Cov}(\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}})$. Define $\textbf{1}_{(j,k)}=\textbf{1}(X_{i}=e_{j}(m_{1})e_{k}(m_{2})^{T})$ (for simplicity of notation, we do not include a subscript i) and denote $I(j,k)=\epsilon_{i}\textbf{1}_{(j,k)}\textbf{1}_{\chi_{i}}$. Note the following facts
$\mathbb{E}[I(j,k)]=\mathbb{E}[\epsilon_{i}]\mathbb{E}[\textbf{1}_{(j,k)}\textbf{1}_{\chi_{i}}]=0,$
$I(j,k)I(l,h)=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).$
Thus $\text{Cov}\big{(}I(j,k),I(l,h)\big{)}=\mathbb{E}[I(j,k)I(l,h)]-\mathbb{E}[I(j,k)]E[I(l,h)]=0$ for $(j,k)\neq(l,h)$. Then we obtain
$\text{Cov}(\bar{\xi}_{i}X_{i})\Big{(}(j,k),(l,h)\Big{)}=\begin{cases}\mathbb{P}(\chi_{i}|\mathcal{E}_{i}(j,k))P_{jk}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)=(l,h),\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (j,k)\neq(l,h).\end{cases}$
Since $P_{jk}\leq\frac{L_{3}}{m\log^{3}d}$, $\|\operatorname{Cov}(\bar{\xi}_{i}X_{1})\|\leq\frac{L_{3}}{m\log^{3}d}$. Thus, $g(Q)\leq\sqrt{\frac{L_{3}}{nm\log^{3}d}}$ holds.
4.$R(Q)\leq\frac{2}{n}$. Obviously, $\|Q_{i}\|=|\textbf{1}_{\chi_{i}}|/n$.
Plug the four estimates for $\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}$ in (28) and we conclude (32) given $n\geq m\log^{4}d$. ∎
4.6 Concentration of Maximum of Independent Random Variables
Given $\|X\|_{\psi_{\alpha}}\leq\sigma$, it holds for $t>0$
$\mathbb{P}(|X|\geq\sigma t)\leq 2\exp(-t^{\alpha}).$
For a metric space $(T,\rho)$, a sequence of its subsets $\{T_{m}\}_{m\geq 0}$ is called an admissible sequence if $|T_{0}|=1$ and $|T_{m}|\leq 2^{2^{m}}$. For $0<\alpha<\infty$, the functional $\gamma_{\alpha}(T,d)$ is defined by
$\gamma_{\alpha}(T,\rho)=\inf_{\{T_{n}\}^{\infty}_{n=1}}\sup_{t\in T}\sum_{n=0}^{\infty}2^{n/\alpha}\rho(t,T_{n})$
where the infimum is taken over all possible admissible sequences. If a random process $\{X_{t},t\in T\}$ satisfies
$\mathbb{P}(|X_{t}-X_{s}|\geq t\rho(t,s))\leq 2\exp(-t^{\alpha}),$
the following theorem holds.
Theorem 15 (Theorem 3.2 in Dirksen [2015]).
There exist constant $C_{1,\alpha},C_{2,\alpha}$ such that
$\mathbb{P}\left(\sup_{t\in T}|X_{t}-X_{t_{0}}|\geq e^{1/\alpha}\left(C_{1,\alpha}\gamma_{\alpha}(T,\rho)+tC_{2,\alpha}\Delta_{\rho}(T)\right)\right)\leq\exp\left(-t^{\alpha}/\alpha\right)$ (34)
where $\Delta_{\rho}(T)=\sup_{t,s\in T}\rho(t,s)$ and $t\geq 0$.
Let $\{\xi_{i}\}_{i=1}^{n}$ be a sequence of independent $\psi_{\alpha}$ random variables with $\|\xi_{i}\|_{\psi_{\alpha}}\leq\sigma$. By Lemma A.3 in Götze et al. [2021], we have the following triangle inequality,
$\|X+Y\|_{\psi_{\alpha}}\leq K_{\alpha}(\|X\|_{\psi_{\alpha}}+\|Y\|_{\psi_{\alpha}}),$
where $K_{\alpha}=2^{1/\alpha}$ for $\alpha\in(0,1)$ and $K_{\alpha}=1$ for $\alpha\geq 1$. We define the trivial metric on $T=\{1,2,\cdots,n\}$,
$\rho(s,t)=\begin{cases}2K_{\alpha}\sigma,t\neq s,\\ 0,t=s.\end{cases}$
With this metric, we construct an admissible sequence $\{T_{n}\}$. Note that there exists $\tilde{m}$ such that $2^{2^{\tilde{m}}}\leq n\leq 2^{2^{\tilde{m}+1}}$. For $m\leq\tilde{m}+1$, take $T_{m}=\{1\}$. For $m>\tilde{m}+1$, take $T_{m}=T$. Then, we obtain
$\sup_{t\in T}\sum_{m=0}^{\infty}\rho(t,T_{n})=\sup_{t\in T}\sum_{m=0}^{\tilde{m}+1}2^{m/\alpha}\rho(t,T_{m})=2K_{\alpha}\sigma\sum_{m=0}^{\tilde{m}+1}2^{m/\alpha}=\sigma C_{\alpha}2^{\tilde{m}/\alpha},$
where $C_{\alpha}$ is a constant only dependent on $\alpha$. Noting that $2^{\tilde{m}}\apprle\log n$, we obtain
$\gamma_{\alpha}(T,\rho)\leq C_{\alpha}\sigma(\log n)^{1/\alpha}.$
Note
$\sup_{i\in\{1,\cdots,n\}}|\xi_{i}|\leq|\xi_{1}|+\sup_{i\in\{1,\cdots,n\}}|\xi_{i}-\xi_{1}|.$
By (34) we get
$\mathbb{P}\left(\sup_{i\in\{1,\cdots,n\}}|\xi_{i}|\geq e^{1/\alpha}\left(C_{1,\alpha}\sigma(\log n)^{1/\alpha}+(C_{2,\alpha}+1)\sigma t\right)\right)$
$\leq$ $\mathbb{P}\left(\sup_{i\in\{1,\cdots,n\}}|\xi_{i}-\xi_{1}|\geq e^{1/\alpha}\left(C_{1,\alpha}\sigma(\log n)^{1/\alpha}+C_{2,\alpha}\sigma t\right)\right)+\mathbb{P}\left(|\xi_{1}|\geq e^{1/\alpha}\sigma t\right)$
$\leq$ $\exp(-t^{\alpha}/\alpha)+\frac{2}{e}\exp(-t^{\alpha}/\alpha)\leq 2\exp(-t^{\alpha}/\alpha),$
where we use $\mathbb{P}(X+Y\geq a+b)\leq\mathbb{P}(X\geq a)+\mathbb{P}(Y\geq b)$. In a cleaner form, we present the concentration inequality
$\mathbb{P}\left(\sup_{i\in\{1,\cdots,n\}}|\xi_{i}|\geq C_{1,\alpha}\sigma(\log n)^{1/\alpha}+C_{2,\alpha}\sigma t^{1/\alpha}\right)\leq 2\exp(-t)$ (35)
As a consequence, if $\{\xi_{i}\}_{i=1}^{n}$ is a sequence of sub-exponential random variables satisfying $\|\xi_{i}\|_{\psi_{2}}\leq\sigma$, there exist constants $C_{1},C_{2}$ such that
$\mathbb{P}\left(\sup_{i\in\{1,\ldots,n\}}|\xi_{i}|\geq C_{1}\sigma\sqrt{\log n}+C_{2}\sigma\sqrt{t}\right)\leq 2\exp(-t).$ (36)
Lemma 16.
Let $Y$ be a random variable with mean 0 and $\mathbb{E}[Y^{2}]\leq\sigma^{2}$. For any $y$,
$\operatorname{Var}(\phi_{y}(Y))\leq\sigma^{2}.$ (37)
Proof.
Let $Y^{\prime}$ be an independent copy of $Y$. Since the function $\psi_{y}(\cdot)$ is Lipchitz with Lipchitz constant $1$,
$2\operatorname{Var}(\psi_{y}(Y))\leq\mathbb{E}(\psi_{y}(Y)-\psi_{y}(Y^{\prime}))^{2}\leq\mathbb{E}(Y^{\prime}-Y)^{2}=2\mathbb{E}[Y^{2}].$
Then, we conclude the lemma. ∎